OpenStudy (anonymous):

solve without a calculator: log(x-4)=log(5x+3)

OpenStudy (anonymous):

since they're logs of the same base, the arguements must be the same: (x - 4) = (5x + 3) now solve this linear equation. and be sure to check your answers

OpenStudy (anonymous):

make each side e^(log(x-4)) = e^(5x-3)) which makes it cancel to (x-4)=(5x-3)

OpenStudy (anonymous):

because even if the algebra is correct, there may be an extraneous solution.

OpenStudy (anonymous):

solving for x, x = ???

OpenStudy (anonymous):

so the answer would be -7/4?

OpenStudy (anonymous):

try again..

OpenStudy (anonymous):

x = -1/4 ?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):